Opened 2 years ago

Closed 22 months ago

#617 closed flight processing (fixed)

GB17/26, flight day 166/2017, Albacete

Reported by: asm Owned by:
Priority: immediate Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/2017/flight_data/spain/RG17_26-2017_166_Albacete

Data arrived from NERC-ARF via hard drive 2017-06-30

Scientific objective: NERC Summer School

Priority: Urgent

PI: Alasdair MacArthur

Sensors:

  • Fenix
  • Leica LIDAR
  • OWL
  • PhaseOne

Change History (31)

comment:1 Changed 2 years ago by asm

Unpacking

Project has been unpacked. There are 8 fenix flightlines and 8 owl lines.

Checks have been run: There are 1 dropped frame incident in 2 different owl lines. Owl nav 4 and owl nav 8 have a few less timestamps than usual.

comment:2 Changed 2 years ago by dac

Navigation

Started navigation processing to test PhaseOne processing chain. No logsheets with basestation so have set height to 0, will need updating and reprocessing later:

lat 39 03 37.39473
long -2 05 53.42671
height 750.819

comment:3 Changed 2 years ago by lah

Owl Processing

Owl files processed to level 1b.

comment:4 Changed 2 years ago by lah

Navigation processing

Had to use 2 base stations to get an ambiguity fix of 2, ALBA and TERU.

Precise positions:
ALBA
Lat: 38 58 40.51145
Long: -1 51 23.03395
Height: 751.633

TERU
Lat: 40 21 01.79014
Long: -1 07 27.48371
Height: 956.187

Not much improvement by altering config settings. Final settings were:
Elevation mask: 10
Q3 ~ 99.5%

Had to set IMU boresight angles all to 0 and also the aircraft frame. Also had to interpolate base station data to 0.01s (Alba) and 0.05s (TERU). Cut off much of the start and end as there was a problem in the data at the start which resulted in failed velocity for an epoch.

comment:5 follow-up: Changed 2 years ago by lah

Owl Mapping

Pretty sure that the sct fix (sct = 2.0) has worked for this flight. Overlapping lines in the same direction 3 & 5 and 2 & 8 align very well. Lines 4 & 6 are pretty well aligned too. However, the opposite direction lines 1 & 7 are slightly offset, which could be because the boresight pitch value is not quite right. Looking into this.

comment:6 in reply to: ↑ 5 Changed 2 years ago by lah

Replying to lah:

Owl Mapping

Pretty sure that the sct fix (sct = 2.0) has worked for this flight. Overlapping lines in the same direction 3 & 5 and 2 & 8 align very well. Lines 4 & 6 are pretty well aligned too. However, the opposite direction lines 1 & 7 are slightly offset, which could be because the boresight pitch value is not quite right. Looking into this.

Have checked various pitch values and 0.33 gives a good alignment of lines 1 and 7. However, this changes the alignment of all the other lines, so happy to stick with the original boresight value of 0.08.

comment:7 Changed 2 years ago by lah

Owl mapping

Sct of 2.12 for line 1 compensates for most of the misalignment. Keeping the other lines as 2.00. Not really any straight edges to determine a better sct if there is one. Proceeding to map final files.

comment:8 Changed 2 years ago by dac

Navigation

Gary provided the height of the basestation from floor to marker (where tape measure slots in) as approximately 1.24 m (from memory). The distance from this to antenna centre is 0.36 m (from previous log sheet) so antenna height to be used for processing is 1.60 m.

comment:9 Changed 2 years ago by dac

Digital camera processing

Converted raw PhaseOne data to tif. Incorrect exposure settings were used so all images are incorrectly exposed but with some post processing might still be usable - at the very least to show the targets laid out for the flight.

comment:10 Changed 2 years ago by dac

Navigation

Reprocessed basestation using antenna height of 1.6 m

lat 39 03 37.39473
long -2 05 53.42670
height 749.115

comment:11 Changed 2 years ago by dac

Navigation

Navigation data processed with basestation only, using provided height.

comment:12 Changed 2 years ago by dac

Hyperspectral Processing

Found SCT values, with Specim fix:

Flightline FENIX
1 2
2 2
3 2
4 1
5 2
6 2
7 2
8 1

comment:13 Changed 2 years ago by dac

Hyperspectral Processing

Ran through mapping on all bands. Will rerun aplxml and create delivery once 2017 data quality report is ready.

comment:14 Changed 2 years ago by lah

Owl processing

Have mapped files with new nav file and created delivery. Ran into a problem creating screenshots, but this was resolved by specifying the projection.

Ready for delivery check. Note that there are no badpixelmethod files for owl.

comment:15 Changed 2 years ago by dac

Digital camera processing

Tagged tiff images, ready for delivery generation.

comment:16 Changed 2 years ago by asm

OWL Delivery Check

Everything looks fine. Will zip mapped files and should be ready to go.

comment:17 Changed 2 years ago by asm

OWL Delivery Check

Files have been mapped. Will mark the project as ready to be delivered.

comment:18 Changed 2 years ago by dac

Hyperspectral Processing

Delivery created - waiting checking.

comment:19 Changed 2 years ago by lah

Hyperspectral Delivery Check

The other links are a bit confused as well, there's no description for EOS and the Report looks like it is the same link as the ASTER page.

  • There's no dem in the delivery!
  • Logsheet - flight overview isn't very detailed (it is just the name). Should we delete the lidar pages this year since the lidar wasn't working?
  • The FOV vectors are for our Fenix, not the loan system.
  • The sensor ID in all the files is for our Fenix, not the loan system.
  • Sensor specifics and sensor ID in flight line xml files incorrect. Our system, not loan.
  • can't run check_apl_cmd without dem
  • Looks like many underflows are not being picked up. Line 1 band 414 - 437, 493 - 545 (507, 514-522 are in readme) Definitely needs to be looked at. Line 2 and 3 have similar problem, probably will apply to other lines.

Many of these issues will probably apply to other projects, so need to be addressed asap. If the wrong FOV vectors have been used to process, then data may need to be reprocessed.

comment:20 Changed 2 years ago by dac

Hyperspectral processing

Started reprocessing with new correct FOV file and updated boresight values.

comment:21 Changed 2 years ago by dac

Hyperspectral processing

Generated new delivery file - ready for checking.

comment:22 Changed 2 years ago by lah

Fenix Delivery Check

  • Few remaining bad pixels in noisy water absorption bands again


Everything else looks ok so marking as ready to deliver and zipping mapped files.

comment:23 Changed 2 years ago by dac

Delivery

Sent data to PI via FTP (slot 5).

comment:24 Changed 2 years ago by asm

Digital Camera DC

Dan created delivery as part of the new processing chain implementation, I took over the DC:

-Google Earth displays the kml fine (and can click in photographs) and the eventfile also looks good on qgis
-No quality reports yet
-Pics are pretty blue (I think this is the flight with wrong camera settings, first flight with new camera). Need to note this on the readme.
-1 Pic is outside flightlines coverage (num 140, but its position is right when checked on google maps)

comment:25 Changed 2 years ago by dac

Digital camera

Added note on blue tinge to images. Will add data quality report once it has been finalised.

comment:26 Changed 2 years ago by dac

Digital camera

Added data quality report. After checking everything has been addressed from DC, ready for delivery.

comment:27 Changed 2 years ago by asm

Digital Camera DC

Everything looks fine, noted sentence about incorrect settings in the readme file. Photograph 140 is outside flightlines coverage but the location is correct so will mark the project as ready to be delivered.

comment:28 Changed 2 years ago by dac

Digital Camera

Delivered via FTP (slot 5). Sent notification to PI.

comment:29 Changed 22 months ago by dac

Archiving

Tidied up and started uploading to CEDA.

comment:30 Changed 22 months ago by dac

Archiving

Data uploaded to CEDA.

comment:31 Changed 22 months ago by dac

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.